# An egonet-based approach to effective weighted network comparison

### v2

@@ Methods
Figure 7: A pictorial representation of an egonet, composed of the ego-node $i$ (center) and its neighbors (red nodes): different edge thickness denotes different weight. The weighted degree $d_{i}$ depends only on the connectivity of node $i$ (red edges); the clustering coefficient $c_{i}$ also depends on the connectivity between the neighbors of $i$ (blue edges); the egonet persistence $p_{i}$ captures the balance between internal and external connectivity (yellow edges).
@@
The performance of each distance is evaluated within the usual Precision-Recall framework: two networks form an actual positive pair if they are generated by the same model, an actual negative pair otherwise. Given a distance $D_{x}$ and a threshold $\varepsilon>0$, a pair of networks is a predicted positive sample if $D_{x}<\varepsilon$, a predicted negative sample otherwise. At this point Precision and Recall are given, for each $\varepsilon$, by $P_{\varepsilon}=TP/(TP+FP)$ and $R_{\varepsilon}=TP/(TP+FN)$, where TP, FP, and FN are, respectively, the number of true positive, false positive, and false negative network pairs. The Precision-Recall curve provides a graphical representation of the simultaneous evolution of $P$ and $R$ with $\varepsilon$, and the Area Under the Precision-Recall curve, $0\leq\text{AUPR}\leq 1$, summarizes the performance of each distance, with the limit $\text{AUPR}=1$ obtained in the ideal case[44, 45].
@@
Given a network of $N$ nodes, we can assume that reading the weight $w_{ij}$ of a connection $(i,j)$ takes a fixed time, so computing the weighted degree of a node takes time $O(N)$, which becomes $O(N^{2})$ for all nodes. For the clustering coefficient (eq. (4)), reading the weights of all connections between neighbors of a node of degree $m_{i}$ requires $O(m_{i}^{2})$ operations, which is $O(N^{2})$ in the worst case and therefore $O(N^{3})$ for the entire network. Similarly for the egonet persistence (eq. (5)): for the numerator we need the weights of the $(m_{i}+1)^{2}$ possible internal connections to $E_{i}$, while in the denominator we read the $(m_{i}+1)N$ possible connections of the nodes of $E_{i}$ with all the nodes of the network: both terms are $O(N^{2})$ in the worst case, which leads to a complexity $O(N^{3})$ for all $N$ nodes. Typical networks are often sparse and therefore quite far from the worst case: we expect lighter computational requirements, but still in the range $O(N^{2})$ to $O(N^{3})$. Indeed, empirical analyzes on the same heterogeneous set of synthetic networks described above obtain a computation time approximately increasing as $t\propto N^{\alpha}$, with $\alpha$ ranging between $2.32$ and $2.50$ for the distances which uses all the three egonet features (see Supplementary Information for details). On the other hand, the worst-case computational requirement of the benchmark distances used for comparison varies from $O(LN+N^{2}\log N)$ for Portrait Divergence on weighted networks[34] (which is $O(N^{3})$ in the worst case $L\approx N^{2}$ and $O(N^{2}\log N)$ for sparse networks $L\approx N$), to $O(N^{3})$ for spectral methods that require the computation of $N$ eigenvalues[55].
@@ Acknowledgements
C.P. was partially supported by the project “CODE – Coupling Opinion Dynamics with Epidemics”, funded under PNRR Mission 4 ”Education and Research” - Component C2 - Investment 1.1 - Next Generation EU ”Fund for National Research Program and Projects of Significant National Interest” PRIN 2022 PNRR, grant code P2022AKRZ9, CUP B53D23026080001.

### v3

@@ Methods
Figure 7: A pictorial representation of an egonet, composed of the ego-node $i$ (center) and its neighbors (red nodes): different edge thickness denotes different weight. The weighted degree $d_{i}$ depends only on the connectivity of node $i$ (red edges); the clustering coefficient $c_{i}$ also depends on the connectivity between the neighbors of $i$ (red and blue edges); the egonet persistence $p_{i}$ captures the balance between internal and external connectivity (red, blue, and yellow edges).
@@
  • $D_{WD}$ (WD-metric): It is defined as a linear combination of three distance terms[34, 35]: two of them compare, with appropriate metrics, the distributions of the (weighted) shortest-path lengths in the two graphs, while the third compares the distributions of the alpha-centralities in the two graphs and their complements.
@@
The evaluation of each distance is carried out within the standard Precision-Recall framework. Two networks are classified as an actual positive pair if they originate from the same model, and as an actual negative pair if they do not. Given a distance measure $D_{x}$ and a threshold $\varepsilon>0$, a pair of networks is considered a predicted positive sample if $D_{x}<\varepsilon$, and a predicted negative sample otherwise. Precision and Recall for each $\varepsilon$ are then defined as $P_{\varepsilon}=TP/(TP+FP)$ and $R_{\varepsilon}=TP/(TP+FN)$, where TP, FP, and FN represent the counts of true positive, false positive, and false negative network pairs, respectively. The Precision-Recall curve visually captures the joint variation of $P$ and $R$ as $\varepsilon$ changes, and the Area Under the Precision-Recall curve (AUPR), with $0\leq\text{AUPR}\leq 1$, encapsulates the performance of the distance measure. The ideal case corresponds to an AUPR of 1[46, 47].
@@
Consider a network with $N$ nodes. The time required to read the weight $w_{ij}$ of a connection $(i,j)$ is constant, meaning that computing the weighted degree of a single node takes $O(N)$ time, and for all nodes, this results in $O(N^{2})$ time. For calculating the clustering coefficient (eq. (4)), determining the weights of all connections between neighbors of a node with degree $m_{i}$ requires $O(m_{i}^{2})$ operations. In the worst case, this is $O(N^{2})$, and for the entire network, the complexity is $O(N^{3})$. Similarly, for the egonet persistence (eq. (5)), the numerator involves reading the weights of the $(m_{i}+1)^{2}$ possible internal connections to $E_{i}$, while the denominator requires reading the $(m_{i}+1)N$ connections between the nodes of $E_{i}$ and all other nodes in the network. In the worst case, both terms are $O(N^{2})$, resulting in an overall complexity of $O(N^{3})$ for the entire network. Typical networks are often sparse and therefore quite far from the worst case: we expect lighter computational requirements, but still in the range $O(N^{2})$ to $O(N^{3})$. Indeed, empirical analyzes on the same heterogeneous set of synthetic networks described above obtain a computation time approximately increasing as $t\propto N^{\alpha}$, with $\alpha$ ranging between $2.32$ and $2.50$ for the distances which uses all the three egonet features (see Supplementary Information for details). On the other hand, the worst-case computational requirement of the benchmark distances used for comparison varies from $O(LN+N^{2}\log N)$ for Portrait Divergence on weighted networks[36] (which is $O(N^{3})$ in the worst case $L\approx N^{2}$ and $O(N^{2}\log N)$ for sparse networks $L\approx N$), to $O(N^{3})$ for spectral methods that require the computation of $N$ eigenvalues[57].
@@ Acknowledgements
C.P. was partially supported by the project “CODE – Coupling Opinion Dynamics with Epidemics”, funded under PNRR Mission 4 ”Education and Research” - Component C2 - Investment 1.1 - Next Generation EU ”Fund for National Research Program and Projects of Significant National Interest” PRIN 2022 PNRR, grant code P2022AKRZ9, CUP B53D23026080001. The author is grateful to the two anonymous Reviewers for their helpful and constructive comments.